# Step1X-Edit: A Practical Framework for General Image Editing

### v3

@@ 4.1 GEdit(Genuine Edit)-Bench
In order to assess the performance of image editing models comprehensively, we have curated a novel benchmark named GEdit-Bench. The core impetus behind developing this benchmark lies in gathering real-world user image editing instances. This enables us to thoroughly examine whether existing editing algorithms can effectively meet practical user editing requirements. Specifically, we have collected over 1,000 user editing examples from platforms like Reddit across the Internet. Subsequently, these editing instructions have been carefully categorized into 11 distinct groups manually. To ensure the diversity of the proposed benchmark, we systematically filtered out instructions with overlapping intents. Finally, we obtain a set of 606 testing samples, all featuring reference images sourced from authentic real-life scenarios. This characteristic renders the benchmark highly applicable for practical use cases.
@@
Instruct-Pix2Pix [8] 3.473 5.601 3.631 4.836 6.948 4.655 3.575 5.491 3.684 4.772 6.870 4.576
MagicBrush [66] 4.646 5.800 4.578 5.806 7.162 5.632 4.677 5.656 4.518 5.733 7.066 5.536
AnyEdit [64] 3.177 5.856 3.231 3.583 6.751 3.498 3.178 5.820 3.212 3.438 6.729 3.361
OmniGen [61] 6.070 5.885 5.162 7.022 6.853 6.565 5.963 5.888 5.061 6.900 6.781 6.413
Step1X-Edit 7.183 6.818 6.813 7.380 7.229 7.161 7.091 6.763 6.701 7.332 7.204 7.104
Gemini [15] 6.697 6.638 6.322 7.276 7.306 6.978 6.732 6.606 6.315 7.287 7.315 6.982
Doubao [50] 7.004 7.215 6.828 7.417 7.635 7.273 6.916 7.188 6.754 7.382 7.639 7.241
GPT-4o [37] 7.844 7.592 7.517 7.873 7.690 7.694 7.850 7.620 7.534 7.826 7.689 7.646
@@
Gemini [15] 5.580 6.757 5.505 5.658 7.362 5.522 5.427 6.767 5.360 5.617 7.360 5.485
Doubao [50] 7.076 7.315 6.869 7.122 7.669 7.062 6.984 7.273 6.772 7.116 7.667 7.063
GPT-4o [37] 7.722 7.590 7.353 7.771 7.625 7.572 7.673 7.559 7.302 7.698 7.634 7.529
Step1X-Edit 7.250 6.855 6.898 7.347 7.327 7.232 7.204 6.869 6.861 7.282 7.303 7.161
@@ 4.2.1 Evaluation on GEdit-Bench
Based on the GEdit-Bench, we evaluated a diverse range of image editing algorithms, covering state-of-the-art open-source solutions such as Instruct-Pix2Pix [8], MagicBrush [66], AnyEdit [64], OmniGen [61], as well as proprietary algorithms like GPT-4.1 [37] [^1] The results are obtained based on ChatGPT APP in April 2025., doubao [50] [^2] The results are obtained based on doubao APP in April 2025., and Gemini2 Flash [15] [^3] The results are obtained in April 2025.. Following VIEScore [25], we adopt three metrics: SQ (Semantic Consistency), PQ (Perceptual Quality), and O (Overall Score). SQ assesses the degree to which the edited results conform to the given editing instruction, with a score ranging from 00 to $10$. PQ evaluates the naturalness of the image and the presence of artifacts, also using a scoring scale that ranges from 00 to $10$. The overall score is calculated based on these evaluations. To perform the automatic evaluation for VIEScore, we adopt the state-of-art MLLM model GPT-4o [^4] API access as of April 2025. Also, the evaluation based on the open-source model Qwen2.5-VL-72B [3] is included for reproduction. To comprehensively assess model capabilities on different languages, each image in our benchmark is paired with one English (EN) and one Chinese (CN) instruction. For EN instructions (GEdit-Bench-EN), both closed and open-source models are evaluated. For CN instructions (GEdit-Bench-CN), only those models which supports Chinese prompts, i.e., the close-source systems, are tested. During the evaluation process, we find that close-source image editing system such as GPT-4o may refuse certain instructions due to safety policies. To address this issue, we report two scores for two testing sets: (1) Intersection Subset — the subset of images whose results can be successfully returned from all the tested models, and (2) Full set — all the testing samples from GEdit-Bench. For the full set of results, we will calculate the average scores only for the cases where the models successfully generate and return the target image. For each evaluated model, instances where no result image is returned due to reasons such as safety concerns will be excluded from the averaging process.

### v4

@@ 4.1 GEdit(Genuine Edit)-Bench
Instruct-Pix2Pix [8] 3.335 6.210 3.234 4.833 6.992 4.691 3.296 6.189 3.219 4.746 6.913 4.578
MagicBrush [66] 4.564 6.335 4.236 5.814 7.149 5.653 4.517 6.371 4.185 5.752 7.069 5.558
AnyEdit [64] 3.122 5.865 2.919 3.873 6.754 3.789 3.053 5.882 2.854 3.713 6.730 3.635
OmniGen [61] 6.037 5.856 5.154 7.033 6.775 6.557 5.879 5.871 5.005 6.845 6.700 6.352
Step1X-Edit 7.289 6.962 6.618 7.501 7.264 7.189 7.131 6.998 6.444 7.388 7.279 7.067
Gemini [15] 6.816 7.408 6.483 7.295 7.314 6.996 6.866 7.436 6.509 7.274 7.327 6.971
Doubao [50] 7.396 7.899 7.137 7.427 7.651 7.285 7.222 7.885 6.983 7.353 7.651 7.230
GPT-4o [37] 7.867 8.097 7.590 7.905 7.723 7.752 7.743 8.133 7.494 7.847 7.705 7.692
@@
Gemini [15] 5.316 7.571 5.208 5.658 7.372 5.566 5.259 7.600 5.143 5.622 7.370 5.525
Doubao [50] 7.228 7.800 6.915 7.109 7.687 7.054 7.168 7.794 6.837 7.098 7.676 7.040
GPT-4o [37] 7.606 7.991 7.336 7.772 7.658 7.599 7.518 8.023 7.301 7.726 7.652 7.552
Step1X-Edit 7.464 7.076 6.779 7.527 7.410 7.259 7.299 7.142 6.658 7.490 7.384 7.212
@@ 4.2.1 Evaluation on GEdit-Bench
Based on the GEdit-Bench, we evaluated a diverse range of image editing algorithms, covering state-of-the-art open-source solutions such as Instruct-Pix2Pix [8], MagicBrush [66], AnyEdit [64], OmniGen [61], as well as proprietary algorithms like GPT-4.1 [37] [^1] The results are obtained based on ChatGPT APP in April 2025., doubao [50] [^2] The results are obtained based on doubao APP in April 2025., and Gemini2 Flash [15] [^3] The results are obtained in April 2025.. Following VIEScore [25], we adopt three metrics: SQ (Semantic Consistency), PQ (Perceptual Quality), and O (Overall Score). SQ assesses the degree to which the edited results conform to the given editing instruction, with a score ranging from 00 to $10$. PQ evaluates the naturalness of the image and the presence of artifacts, also using a scoring scale that ranges from 00 to $10$. The overall score is calculated based on these evaluations. To perform the automatic evaluation for VIEScore, we adopt the state-of-art MLLM model GPT-4o [^4] API access as of June 2025. Also, the evaluation based on the open-source model Qwen2.5-VL-72B [3] is included for reproduction. To comprehensively assess model capabilities on different languages, each image in our benchmark is paired with one English (EN) and one Chinese (CN) instruction. For EN instructions (GEdit-Bench-EN), both closed and open-source models are evaluated. For CN instructions (GEdit-Bench-CN), only those models which supports Chinese prompts, i.e., the close-source systems, are tested. During the evaluation process, we find that close-source image editing system such as GPT-4o may refuse certain instructions due to safety policies. To address this issue, we report two scores for two testing sets: (1) Intersection Subset — the subset of images whose results can be successfully returned from all the tested models, and (2) Full set — all the testing samples from GEdit-Bench. For the full set of results, we will calculate the average scores only for the cases where the models successfully generate and return the target image. For each evaluated model, instances where no result image is returned due to reasons such as safety concerns will be excluded from the averaging process.